# GroundGazer: Camera-based indoor localization of mobile robots with millimeter accuracy at low cost

### v3

@@ III-B QR code detection and masking
QR codes are detected using OpenCV, which outputs the four corner points of the QR code in known order. This gives the global orientation information, previously shown in Fig. 1(b). The QR code ID is extracted with the pyzbar library [^6] Availabe at: https://pypi.org/project/pyzbar/, as it gave better results than OpenCV. The QR ID enables identification of the respective chessboard square. By computing the difference in $x$ and $y$ between the square centers where the QR code is detected and where the crosshair is detected, the QR ID also gives us the square in which the crosshair lies and thus global localization capabilities.
Since the QR codes would disturb the successive image processing operations, we mask them out. Therefore, the average color around the detected QR code is computed and the region of the QR code is replaced with a slightly larger rectangle of this color. This masking procedure might also mask out parts of the crosshair if it lies on a QR code. The QR code masking could be improved in future work, but our experiments have shown that our system can cope with irregular missed crosshair detections. By proper parametrization of the Hough transform in the crosshair detection, we can also fully reconstruct a partly occluded crosshair.
III-C Crosshair detection
The green crosshair can be extracted by filtering the image with a mask that keeps the green image parts. This mask has to be carefully tuned as both the different colors of the chessboard and varying illumination condition influence the detected crosshair color. We therefore apply a concatenation of multiple masks. The first mask filters out all pixels with low green values. We then convert the image to the HSV color space, which was more robust than only working in the RBG space. Three additional masks filter out pixels with low saturation, low hue, and pixels where the sum of hue and value is low. A combined crosshair mask is found in Fig. 3(a). The horizontal and vertical lines in a region of 150 pixels around estimated crosshair position from the previous time step are then found by Hough transformation [hough_paper], clustered into lines with similar angle and start- and endpoints, and the start and end $x$ values of the clustered lines are averaged for the vertical, and the $y$ values for the horizontal line. The crosshair location in image coordinates is then estimated as the intersection of the horizontal and vertical crosshair lines. The final crosshair lines and the detected crosshair position are shown in Fig. 3(b).
@@ III-C Crosshair detection
If the crosshair is not detected or if there is a large deviation from the expected crosshair position, we use an average of the previous crosshair positions as estimate. In principle, our system can work without the laser diode by calibration one initial crosshair and using a virtual crosshair afterwards, which stays in the same position in the image under ideal conditions. In the experiments with our low-cost robot that has limited mechanical stability, we have found that using the laser diode, which is lightweight and mounted much closer to the ground than the camera, is more accurate and robust than using a virtual crosshair due to less vibrations and shaking of the laser diode.
@@ III-D Chessboard square detection
To detect the chessboard squares, we first apply a Canny filter for edge detection after grayscale conversion and Gaussian smoothing and mask out the previously detected laser crosshair (after extending the crosshair lines), as shown in Fig. 4(b). The binary Canny output is then processed by Hough transformation to find the chessboard lines. For further denoising, the detected lines are broadened and the image of lines is processed by an opening operation, which consists of erosion followed by dilation. In dilation and erosion, a 2D kernel is slided over the image, where the kernel center corresponds to the pixel to process. In each kernel, erosion only keeps white (detected line) pixels if all kernel pixels are white. Dilation does the opposite and sets the color of the center pixel to white if any pixel in the kernel is white. The image after Hough transformation and opening is depicted in Fig. 4(c). For further operations, the image is inverted, as seen in Fig. 4(d). The chessboard squares are detected by first finding and simplifying the white square contours with the functions findContours [find_contours] and approxPolyDP [ramer, douglas_peuker]. The simplified contours (polygons) are then filtered for square detection. Only polygons that have four edges, relative side lengths between factor 0.95 and 1.05, and an area of 80.000 - 100.000 pixels are kept. As the detected squares are slightly smaller due to the line broadening and opening, 10 pixels are added to the square corner coordinates. Those detected squares are given in Fig. 4(e). In the last step, the detected corners are refined to subpixel accuracy with cornerSubPix [subpix] shown in Fig. 4(f).
@@
((a)) Top-down view.
((b)) Canny edges after crosshair and QR code masking.
@@ III-E Position and heading estimation
Given the detected corner points of the chessboard square with detected laser crosshair, we compute the relative position of the crosshair in the square. We then convert the crosshair pixel coordinates to global coordinates with the known square size of $16.\overline{6}$ cm. To resolve the rotation ambiguity of the local square coordinate system and compute global world coordinates, the QR code references are used, as previously shown in Fig. 1(b). The heading is computed as the angle between the chessboard lines in $x$ direction and the vertical crosshair line. We currently get the chessboard lines though a line filter, but future work should simply use the chessboard corners with sub-pixel accuracy for this task.
@@
If no QR code is detected as rotation reference, we predict the current heading using a constant velocity model. The heading is then estimated by choosing the detected angle from the set of ambiguous angles (occuring due to the lack of knowledge which chessboard lines point in $x$- and $y$-direction) by a nearest neighbor assignment. Since the AMR has to move in a physically plausible way and the heading cannot make large jumps between two images, this works reliably.
@@
AMR position and heading estimation is also possible without QR codes through crosshair tracking in a simple fashion. Since we can reliably estimate the heading, no position ambiguity in local square coordinates occurs while tracking the crosshair as the order of P1-P4 in Fig. 1(b), i.e. the orientation of a locally observed square, stays known. To catch jumps between squares, we check if the crosshair position in local square image coordinates in $x$ or $y$ surpasses a threshold of $150$ pixels and if so, we perform a global square jump in the respective direction. [^7] A preliminary version of GG that didn’t contain QR codes yet gave similar position and heading estimation results as with QR codes, but this allows only relative position estimation w.r.t. the initial position and has no mechanism for position recovery after failure or kidnapping.
@@
Using this crosshair tracking after an initial global position and heading estimate with a QR code, further frequent global updates with the QR code references are not necessary, but this requires a high enough image sampling frequency such that the Nyquist-Shannon sampling theorem holds. The spatial frequency $f_{pattern}$ of the chessboard pattern is $f_{pattern}=\frac{1}{16.\overline{6}}\,\mathrm{cm}^{-1}$ and thus the images have to be sampled with a spatial frequency of at least $f_{image}=2f_{pattern}$ to avoid aliasing (causing jumps into the wrong square while crosshair tracking). This upper bounds the maximum AMR velocity to $v_{max}=f_{image}^{-1}f_{fps}$, where $f_{fps}$ is the sampling rate of the camera. We record with $f_{fps}=8.\overline{3}$ Hz, i.e. $v_{max}\approx 0.69\,\frac{\mathrm{m}}{\mathrm{s}}$, which is more than double of the maximum velocity of our AMR. To adapt the crosshair tracking for faster AMRs, one would have to either increase the square size, the sampling rate of the camera, or add a robust crosshair position prediction model (e.g. through odometry).
@@
III-F Trajectory smoothing
To increase the accuracy of the estimated AMR positions and headings, we smooth them in post-processing. This is especially helpful in case of vibrations and shaking when the AMR moves. Since we want to use the GG system as an offline reference system for another novel, radar-based indoor localization system developed by us and described in [ipin_pascal, ipin_sven, irs25_raytracing], we choose polynomial fitting for this task. An online version, e.g. with an alpha-beta filter or with a Kalman filter [bar-shalom_estimation_2002] would, however, be straightforward.
We split the estimated trajectory into equally sized windows with 50% overlap. We then fit a polynomial of given degree to the respective $x$, $y$ and $\theta$ values in each window. We experimentally set the window size to 20 and the polynomial degree to 5. The smoothed values in the overlapping windows are averaged for better estimates and transitions between windows.
@@ IV-B Localization experiments
To showcase the correct working of our system under real inference conditions with possible motion blur, vibrations etc., we drive the AMR in the laboratory by steering it wirelessly with a PS4 controller. Since hand-collecting ground truth on the floor [ground_truth_laser] (e.g. with a set square and a protractor) is an error prone and tedious task that only works in a static setting, we developed a software tool to label the ground truth position and heading based on the corresponding top-down image by hand.
@@ V Results
Our trajectory results are visualized in Fig. 5.
((a)) Full crosshair trajectories.
((b)) Zoomed-in region.
Figure 5: Trajectory evaluation with a) full trajectories of the crosshair before and after smoothing and b) zoom into trajectory. Smoothing reduces the jitter in the raw trajectory caused by the shaking of the AMR.
The full true and estimated raw and smoothed trajectories are shown in Fig. 5(a). As visible, the estimates closely follow the ground truth. The raw crosshair trajectory shows some jitter. This is caused by the shaking of the low-cost robot and not by the image processing. With the proposed smoothing, the jitter is reduced and the trajectory is improved, as evident from the zoomed-in region in Fig. 5(b).
@@
((b)) Cumulative density function of the absolute errors.
@@
A closer error evaluation is found in Fig. 6. The errors in $x$, $y$ and $\theta$ are depicted in Fig. 6(a). All errors in $x$ and $y$ are below 1 cm. A few outliers arise due to small errors in the crosshair or corner detection, especially when their lines overlap. Heading errors are mostly below 1 degree. Heading has a few more outliers than $x$ and $y$ because the crosshair can get skewed by uneven ground or by the shaking of the AMR. Heading has also proven to be harder to label than $x$ and $y$. It should be noted that we introduce errors in the ground truth with our labeling method and we believe our system to be more accurate. We skip the error evaluation with smoothing, as it only improved the realness of the crosshair trajectory and not the accuracy. Smoothing is useful when mapping from the crosshair to a sensor from a different localization system on the AMR, where jitter degrades the accuracy of the crosshair reference. The corresponding CDFs of the absolute errors are found in Fig. 6(b). They show clearly that most errors in $x$ and $y$ are in low mm or even sub-mm range.
Table. I gives the bias, the mean absolute error (MAE), the maximum error, the 95th percentile and the 99th percentile of the absolute error for each parameter.
Table I: Error statistics for $x$ and $y$ in mm and for $\theta$ in degree.
@@
$x$ -0.02 0.89 2.96 5.04 6.44
$y$ 0.04 0.66 1.95 3.18 4.08
$\theta$ -0.35 0.51 1.25 1.65 2.11
The estimates in $x$ and $y$ have negligible bias. Positioning errors in both dimensions are similar as expected. The errors in $x$ are slightly higher due to more outliers as visible in Fig. 6(a). The 99th percentile of the absolute errors in $x$ and $y$ is far below 1 cm. The heading is underestimated as seen in the bias and the MAE is heavily influenced by this bias. Improving the laser setup, using a stable virtual crosshair, or calibration the heading by de-biasing would improve performance. Still, the 95th percentile of heading errors is just slightly over 1 and a non-negligible part of this error is likely due to the labeling process.
@@ VI Discussion
Our proposed GG system is fast, low-cost, and gives mm accuracy. It is, however, just a prototype and this localization approach brings several advantages and can be further improved upon.
@@
Cost and portability: Our system is low-cost due to the low hardware requirements (laser diode, camera, chessboard, processor). We are confident that the system will also work with even cheaper hardware. The largest cost factor is currently the chessboard floor. However, any grid pattern is sufficient. For rooms without such chessboard floor, one might e.g. build a grid by projection orthogonal laser lines on the floor using a laser level and drawing or taping such line pattern on the floor using the laser lines. A grid made of fiducial markers, e.g. with QR codes as in [QR_drones, QR_docking_submm], is another computationally more expensive and most likely less accurate and robust option than a grid of lines. One may also build a portable version, where pre-built chessboard grid patches can be sticked together. The system can further work with a cheaper camera (the current one already costs less than $100\,\mathdollar$) and since the image processing is relatively simple, one might implement a future (online) version that runs on a field programmable gate array (FPGA).
@@
Accuracy and 3D position and orientation estimation: Due to the high spatial resolution of modern cameras and the closeness of the camera to the position references (chessboard corners) and the crosshair on the ground, we achieve high positioning accuracy. The localization accuracy of GG is also mostly independent of the AMR position in the room, given similar illumination conditions. Performance might be enhanced by multi-camera and multi-laser setups. The localization method could be extended to handle arbitrary detected chessboard corners, opposed to the four square corners in which the crosshair lies. Our current positioning bottleneck (with a moving AMR) is the not the localization algorithm, but the mechanical stability of the AMR, which causes larger shaking and vibration of the camera, and to a smaller extend also of the laser diode and the crosshair. Simple improvements would be to use an AMR with better stability, more rigid camera and laser diode installations, and adding some stabilizers for the camera and more importantly, for the laser diode in case of localization using a real crosshair.
@@
An alternative localization approach that could be studied in future work is Perspective-n-Point (PnP) based localization with a monocular camera, as often used with fiducial markers. Therefore, the known correspondences between the detected 2D chessboard corner points in the image and their known 3D positions in world coordinates enable 3D camera position and orientation estimation without the laser diode.
@@
Global localization: By equipping the floor with QR codes, we achieve global localization capabilities. QR codes were, however, just a readily available solution for this proof-of-concept of the system. Simpler visual encodings that are easier and faster to detect would improve the system, especially for an online version. The encodings can be used to either identify the chessboard squares and their corners (as in our current version), or to identify the different grid lines.
@@
Real-time processing: In addition to an FPGA version, crosshair tracking, simpler global position and orientation references, an alpha-beta filter for AMR trajectory smoothing etc., the image processing still has much potential for improvement. Instead of running the full QR code, crosshair and corner detection for each image as described in Sec. III (which itself has room for improvement, e.g. by initial corner detection through a simpler grid pattern finding method at lower resolution instead of the expensive Hough transformation at full resolution), one might track the global references and the periodic chessboard corners after initial detection and make use of the prior knowledge about the crosshair position in the image. This would require only processing smaller patches of the images, thus speeding up the detection chain.
@@
Robustness: The redundant, periodic chessboard pattern allows our system to work even if large areas of the chessboard pattern are occluded, e.g. by other AMRs, people, dirt etc. Since the crosshair stays roughly in the same position of the image and parts of the crosshair are sufficient to reconstruct the full crosshair, we can also deal with crosshair occlusion. Robustness could be further increased by adding more cameras, laser diodes, and odometry, which would additionally improve accuracy as noted before, but also increase system complexity and cost.
@@
Scalability: Our GG system is scalable to large areas. In addition, the system can be scaled to robot swarms. In a version without the laser diode (virtual crosshair or PnP), there will be no interference between robots except for some floor occlusion, which the system can handle. If all robots use the same laser diode, we can still differentiate the different lasers crosshairs, as the crosshair of each respective robot stays in the same place in the images. In case there might be interference between the crosshairs, one could install lasers of different colors on the robots.

### v4

@@ III-B QR code detection and masking
QR codes are detected using OpenCV, which outputs the four corner points of the QR code in known order. This gives the global orientation information, previously shown in Fig. 1(b). The QR code ID is extracted with the pyzbar library [^6] Availabe at: https://pypi.org/project/pyzbar/., as it gave better results than OpenCV. The QR ID enables identification of the respective chessboard square. By computing the difference in $x$ and $y$ between the square centers where the QR code is detected and where the crosshair is detected, the QR ID also gives us the square in which the crosshair lies and thus global localization capabilities.
Since the QR codes would disturb the successive image processing operations, we mask them out. Therefore, the average color around the detected QR code is computed and the region of the QR code is replaced with a slightly larger rectangle of this color. This masking procedure might also mask out parts of the crosshair if it lies on a QR code. The QR code masking could be improved in future work, but our experiments have shown that our system can cope with irregular missed crosshair detections. By proper parametrization of the Hough transform [10] in the crosshair detection, we can also detect an incomplete crosshair.
III-C Crosshair detection and masking
The green crosshair can be extracted by filtering the image with a mask that keeps the green image parts. This mask has to be carefully tuned, as both the different colors of the chessboard and varying illumination condition influence the detected crosshair color. We therefore apply a concatenation of multiple masks. The first mask filters out all pixels with low green values. We then convert the image to the HSV color space, which was more robust than only working in the RBG space. Three additional masks filter out pixels with low saturation, low hue, and pixels where the sum of hue and value is low. A combined crosshair mask is found in Fig. 3(a). The horizontal and vertical lines in a region of 150 pixels around the estimated crosshair position from the previous time step are then found by Hough transformation, clustered into lines with similar angle and start- and endpoints, and the start and end $x$-values of the clustered lines are averaged for the vertical, and the $y$-values for the horizontal line. The crosshair location in image coordinates is then estimated as the intersection of the horizontal and vertical crosshair lines. The final crosshair lines and the detected crosshair position are shown in Fig. 3(b). To not disturb the chessboard square detection, we mask out the crosshair with an extended and broadened version of the detected crosshair.
@@ III-C Crosshair detection and masking
If the crosshair is not detected or if there is a large deviation from the expected crosshair position, we use an average of the previous crosshair positions as estimate. In principle, our system can work without the laser diode by calibration one initial crosshair and using a virtual crosshair afterwards, which stays in the same position in the image under ideal conditions. Since we estimate the heading using only the $x$-direction chessboard lines and we need no crosshair line directions, this would mean fixating one point in the image and using it as AMR position. In the experiments with our low-cost robot that has limited mechanical stability, we have found that using the laser diode, which is lightweight and mounted much closer to the ground than the camera, is more accurate and robust than using a virtual crosshair due to less vibrations and shaking of the laser diode.
@@ III-D Chessboard square detection
To detect the chessboard squares, we first apply a Canny filter for edge detection after grayscale conversion and Gaussian smoothing, as shown in Fig. 4(b). The binary Canny output is then processed by Hough transformation to find the chessboard lines. For further denoising, the detected lines are broadened and the image of lines is processed by an opening operation, which consists of erosion followed by dilation. In erosion and dilation, a 2D kernel is slided over the image, where the kernel center corresponds to the pixel to process. In each kernel, erosion only keeps white (detected line) pixels if all kernel pixels are white. Dilation does the opposite and sets the color of the center pixel to white if any pixel in the kernel is white. The image after Hough transformation and opening is depicted in Fig. 4(c). For further operations, the image is inverted, as seen in Fig. 4(d). The chessboard squares are detected by first finding and simplifying the white square contours with the functions findContours [31] and approxPolyDP [28, 9]. The simplified contours (polygons) are then filtered for square detection. Only polygons that have four edges, relative side lengths between factor 0.95 and 1.05, and an area of 80.000 - 100.000 pixels are kept. As the detected squares are slightly smaller due to the line broadening and opening, 10 pixels are added to the square corner coordinates. Those detected squares are given in Fig. 4(e). In the last step, the detected corners are refined to subpixel accuracy with cornerSubPix [13] as shown in Fig. 4(f).
@@
((a)) Top-down view after QR code masking.
((b)) Canny edges with masked crosshair.
@@ III-E Position and heading estimation
Given the detected corner points of the chessboard square with detected laser crosshair, we compute the relative position of the crosshair in the square. We then convert the crosshair pixel coordinates to global coordinates with the known square size of $16.\overline{6}$ cm. To resolve the rotation ambiguity of the local square coordinate system and compute global world coordinates, the QR code references are used, as previously shown in Fig. 1(b). The heading is computed as the angle between the chessboard lines in $x$-direction and the calibrated vertical image direction.
@@
If no QR code is detected, we predict the current heading using a constant velocity model. The heading is then estimated by choosing the detected angle from the set of four ambiguous angles (due to the lack of knowledge which chessboard lines point in $x$- and $y$-direction) by a nearest neighbor assignment. The four ambiguous chessboard line directions are computed by clustering and averaging the lines from Hough transformation depicted in Fig. 4(c). Since the AMR has to move in a physically plausible way and the heading cannot make large jumps between two images, this nearest neighbor assignment works reliably.
@@
AMR position estimation is also possible without QR codes through crosshair tracking in a simple fashion. Since we can reliably estimate the heading, no position ambiguity in local square coordinates occurs while tracking the crosshair, as the order of P1-P4 in Fig. 1(b), i.e. the orientation of a locally observed square, stays known. To catch jumps between squares, we check if the crosshair position in local square image coordinates in $x$ or $y$ surpasses a threshold of $150$ pixels and if so, we perform a global square jump in the respective direction.
@@
Using this crosshair tracking after an initial global position and heading estimate with a QR code, further frequent global updates with the QR code references are not necessary, but this requires a high enough image sampling frequency such that the Nyquist-Shannon sampling theorem holds. The spatial frequency $f_{pattern}$ of the chessboard pattern is $f_{pattern}=\frac{1}{16.\overline{6}}\,\mathrm{cm}^{-1}$ and thus the images have to be sampled with a spatial frequency of at least $f_{image}=2f_{pattern}$ to avoid aliasing (causing jumps into the wrong square while crosshair tracking). This upper bounds the maximum AMR velocity to $v_{max}=f_{image}^{-1}f_{fps}$, where $f_{fps}$ is the sampling rate of the camera. We record with $f_{fps}=8.\overline{3}$ Hz, i.e. $v_{max}\approx 0.69\,\frac{\mathrm{m}}{\mathrm{s}}$, which is more than double of the maximum velocity of our AMR. To adapt the crosshair tracking for faster AMRs, one would have to either increase the unambiguous region (e.g. by increasing the square size), increase the sampling rate of the camera, or add a robust and accurate crosshair position prediction model (e.g. through odometry).
@@ IV-B Localization experiments
To showcase the correct working of our system under real inference conditions with possible motion blur, vibrations etc., we drive the AMR in the laboratory by steering it wirelessly with a PS4 controller. Since hand-collecting ground truth on the floor [3] (e.g. with a set square and a protractor) is an error prone and tedious task that only works in a static setting, we developed a software tool to label the ground truth position and heading based on the corresponding top-down image by hand. We therefore zoom in and select the crosshair center and three points of the respective square to build the coordinate system for positioning, including the $x$-direction line for heading estimation. The points are selected to the best of our abilities with pixel accuracy, where one pixel equals roughly $\frac{16.66}{300}\simeq 0.56$ mm (see Fig. 1(b)).
@@ V Results
The labeled and estimated trajectories are shown in Fig. 5. As visible, the estimates closely follow the ground truth.
Figure 5: Trajectory evaluation
@@
((b)) Cumulative density functions of the absolute errors.
@@
A closer error evaluation is found in Fig. 6. The errors in $x$, $y$ and $\theta$ are depicted in Fig. 6(a). A few outliers arise due to small errors in the crosshair or corner detection, often caused by overlaps of the crosshair with the chessboard lines and corners. The corresponding CDFs of the absolute errors are found in Fig. 6(b). They show clearly that most errors in $x$ and $y$ are in low mm or even sub-mm range.
Table. I gives the bias, the mean absolute error (MAE), the maximum absolute error (Max), and the 95th and 99th percentiles of the absolute error for each parameter.
Table I: Error statistics for $x$ and $y$ in mm and for $\theta$ in degrees.
@@
$x$ -0.03 0.92 2.62 5.03 6.44
$y$ 0.02 0.69 1.89 3.02 3.70
$\theta$ -0.05 0.42 1.09 1.43 1.47
All estimates have negligible bias. Positioning errors in both dimensions are similar as expected. The errors in $x$ are higher due to more outliers as visible in Fig. 6(a). The 99th percentile of the absolute errors and the max errors in $x$ and $y$ are well below 1 cm.
Heading estimation also shows good performance with the 95th percentile of absolute heading errors just slightly being over $1^{\circ}$ and a non-negligible part of this error likely being due to the labeling process. Good heading accuracy is important to accurately map from the crosshair center to another position around the crosshair, e.g. to a localization sensor on the AMR.
@@ VI Discussion
Our proposed GG system is low-cost and gives mm accuracy. It is, however, just a prototype and this localization approach brings several advantages and can be further improved upon.
@@
Cost and portability: Our system is low-cost due to the low hardware requirements (camera, chessboard, processor, optional laser diode). We are confident that the system will also work with even cheaper hardware. The largest cost factor is currently the chessboard floor. However, any grid pattern is sufficient. For rooms without such chessboard floor, one might e.g. build a grid by projection orthogonal laser lines on the floor using a laser level and drawing or taping such line pattern on the floor using the laser lines. One may also build a portable version, where pre-built chessboard grid patches can be sticked together. The system can further work with a cheaper camera and since the image processing is relatively simple, one might implement a future (online) version that runs on a field programmable gate array (FPGA).
@@
Accuracy: Due to the high spatial resolution of modern cameras and the closeness of the camera to the position references (chessboard corners) and the crosshair on the ground, we achieve high positioning accuracy. The localization accuracy of GG is also mostly independent of the AMR position in the room, given similar illumination conditions. Performance might be enhanced by multi-camera and multi-laser setups. The localization method could be extended to handle arbitrary detected chessboard corners, opposed to the four square corners in which the crosshair lies. Our current positioning bottleneck (with a moving AMR) is the not the localization algorithm, but the mechanical stability of the AMR, which causes larger shaking and vibration of the camera, and to a smaller extend also of the laser diode and the crosshair. The resulting position and heading jitter is not considered in our evaluation. Simple hardware improvements include an AMR with better stability, more rigid camera and laser diode installations, and stabilizers for the camera and also for the laser diode in case of localization using a real crosshair. On the software side, smoothing the trajectory, e.g. online with an alpha-beta or a Kalman filter [2], or offline with a Savitzky-Golay [29] filter, would reduce this jitter.
@@
3D position and orientation estimation: An alternative localization approach that could be studied in future work is Perspective-n-Point (PnP)-based localization with a monocular camera, as often used with fiducial markers and realized in a conceptionally similar system to ours with a hand-held camera and a small, printed chessboard in [30]. Therefore, the known correspondences between the detected 2D chessboard corner points in the image and their known 3D positions in world coordinates enable 3D camera position and orientation estimation without the laser diode. While potentially less accurate in 2D position estimation, this could minimize the error from projecting the estimated position by GG to the position of another localization sensor on the AMR. The current localization method could theoretically employ a (virtual) crosshair located on the floor below the AMR for positioning. However, this requires installing the camera and additional lighting at the bottom of the AMR chassis and dealing with very small camera-ground distance (motion blur & visibility of full chessboard squares).
@@
Global localization: By equipping the floor with QR codes, we achieve absolute, global localization capabilities. QR codes were, however, just a readily available solution for this proof-of-concept of the system. Simpler visual encodings that are easier and faster to detect would improve the system, especially for an online version. The encodings are used to identify the chessboard square corners, e.g. with the known QR corners as in our current version, or by other means such as making the different grid lines identifiable.
@@
Real-time processing: In addition to the implemented crosshair tracking and the proposed FPGA version, simpler global position and orientation references etc., the image processing still has much potential for improvement. Instead of running the full QR code, crosshair and corner detection for each image as described in Sec. III (which itself has room for improvement, e.g. by applying a simpler method run at lower resolution to find the initial chessboard corners instead of the expensive Hough transformation at full resolution), one might track the global references and the periodic chessboard corners after initial detection and make use of the prior knowledge about the crosshair position in the image. This would require processing only smaller patches of the images, thus speeding up the detection chain.
@@
Robustness: The redundant, periodic chessboard pattern allows our system to work even if large areas of the chessboard are occluded, e.g. by other AMRs, people, dirt etc. Since the crosshair stays roughly in the same position of the image and parts of the crosshair are sufficient for crosshair detection, we can also deal with crosshair occlusion. Robustness could be further increased by adding more cameras, laser diodes, and odometry, which would additionally improve accuracy, but also increase system complexity and cost. Moreover, the crosshair could be replaced by a more redundant laser pattern.
@@
Scalability: Our GG system is scalable to large areas. In addition, the system can be scaled to robot swarms. In a version without the laser diode (virtual crosshair or PnP), there will be no interference between robots except for some floor occlusion, which the system can handle. If all robots use the same laser diode, one can still differentiate the different laser crosshairs, as the crosshair of each respective robot stays in the same place in the images. In case there might be interference between the crosshairs, one could try installing lasers of different colors or patterns on the robots.